Re: Shift report 20000805 16:00-24:00

From: Konstantin Olchanski (olchansk@ux1.phy.bnl.gov)
Date: Sun Aug 06 2000 - 13:28:07 EDT

  • Next message: http account: "Shift report 20000806 08:00-16:00"

    > 
    > ------------------------------------------------------------
    > Date: Sat Aug  5 00:04:09 2000, Shift: 16:00-24:00
    > ------------------------------------------------------------
    > 
    > 3. Detectors
    > 
    >     - T2 is out of DAQ stream due to the readout failure
    > ------------------------------------------------------------
    
    
    This is not a T2 readout failure in the sense that nothing
    has failed or broken down.
    
    The problem started when the T2 readout electronics board in the IR was
    replaced, and the replacement board turned out to be a dud. This
    replacement board was a known good board, used in all the TPC tests in
    the lab, and it is not clear why it does not work in the IR.
    
    
    -- 
    Konstantin Olchanski
    Physics Department, Brookhaven National Laboratory, Long Island, New York
    olchansk@bnl.gov
    



    This archive was generated by hypermail 2b29 : Sun Aug 06 2000 - 13:31:26 EDT